OpenStudy (anonymous):

help please hat is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

x-y=-3 -x+y=3 (-x+y)+(x+3y)=3+5=8 4y=8 y=2 x-2=-3 x=-1

OpenStudy (anonymous):

Methods: -graphing -substitution -elimination

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Find the quotient of 20a^3b+15ab^4-25ab over 5ab??